# Good News for Polynomial Root-finding

### v35

@@ D.2 Definition
$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (D.3)
@@
Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [68, 55, 126, 138].
@@
D.3 Doubling precision and the deflation of a wild factor
Computational (working) precision sufficient in order to control the relative error of an output approximation to a root can be defined by root’s condition number (see the relevant discussion in [14, 21]). This number is not readily available, however, and instead one can recursively double precision of computing, until a fixed output accuracy is verified: first apply a fixed root-finding iterations with the IEEE double precision of 53 bits and then recursively double it as long as this is feasible and enables more accurate approximation of the roots, stopping where the required output accuracy is ensured.
This policy for root-finding can be traced back to [14], where it was proposed for MPSolve. It turned out to be highly efficient and was adopted in [41] for Eigensolve, in [21] for the next version of MPSolve, and in [126, 138] for approximation of all roots by means of Newton’s iterations.
According to recipe (D.2) the computed approximation $z$ is only updated until $p(z)=0$ or, in numerical implementation, until the roots become tame, for a fixed stopping criterion. In the latter case, one should restrict updating to the wild roots (cf. Sec. 1.7) and should deflate the wild factor, if it has small degree. When a root is tamed the minimal working precision is exceeded by less than a factor of 2. Recall from [130] and [132, Thm. 2.7] that working precision does not need to exceed the output precision $b^{\prime}$ by more than a factor of $d$, and so at most $O(\log(db^{\prime}))$ steps of doubling precision are sufficient.
For every working precision the $d$ roots of $p(x)$ are partitioned into $d-w$ tame and $w$ wild roots as above; then one can apply explicit or implicit deflation (cf. (1.14)).
[…]
D.4 Acceleration with Fast Multipole Method (FMM)
Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [48] and by Carrier, Greengard, and Rokhlin [27] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The paper [117] proposes and extensively tests its acceleration to Superfast Multipole Method. The papers [101, 103] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
@@ D.4 Acceleration with Fast Multipole Method (FMM)
$W(x)$ of (D.3) and all $j$ (cf. [15]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
@@
$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (D.4)
Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf. [15])). [^50] Gerschgorin’s theorem (cf. [46]) for the matrix $xI_{d}-A$ defines neighborhoods of its eigenvalues (the roots of $p$), which helps to analyze and to modify the iterations (cf. [21]). Equations (D.4) imply that
@@
$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (D.5)
Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (D.3)) by applying FMM.
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (D.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (D.3) and (D.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [21, equation (17)]:
@@
(i) Use the following customary trick (see [30, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [30] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
@@
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (1.14) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 10.1. Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (1.14)) and can apply FMM to accelerate numerical computation of this sum. To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
D.5 Computation of a disc covering all zeros of a black box polynomial
D.5.1 Introductory comments
@@ D.5.3 Deterministic solution
D.6 $d$ root radii and initialization
The algorithms of [68, 55, 138] approximate all $d$ roots of a polynomial $p$ by applying Newton’s iterations at $f(d)$ equally spaced points on a large circle containing inside all $d$ roots for $f(d)$ of order at least $d\log(d)$. Bounds (10.6) immediately define such a circle for a polynomial given by its coefficients.
MPSolve further elaborates upon initialization. It first applies Bini’s heuristic recipe of [10], reused in [14, 21], for the approximation of the $d$ root radii by involving all coefficients of $p$. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization recipe turns out to be highly efficient, and one is challenged to test whether it can also help accelerate convergence of Newton’s iterations.
The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s. It was first outlined and briefly analyzed in [130] for fast approximation of a single root radius and then extended in [88, Sec. 4] (cf. also [89, Sec. 5]) to approximation of all root radii of a polynomial. As in [10, 14, 21] it uses all $d$ coefficients of $p$. A variation of this algorithm in [59] supports the following theorem.
Theorem D.1.
[59, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
@@ D.6 $d$ root radii and initialization
Combine Thm. D.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [130, Cor. 14.3] and obtain
Corollary D.1.
@@
The algorithm supporting the corollary and said to be Alg. D.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [56], which extends the novel ingenious algorithm of [78] for multipoint polynomial evaluation.
@@
(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects $\mathbb{U}$.
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis (see Fig. 10). [122] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$. According to these tests the algorithm is highly efficient for the approximation of all isolated real roots. Clearly, the well-developed root-finding techniques, e.g., ones using derivatives and higher order derivatives (see [99, 62], and the references therein) can further strengthen these algorithms. By complementing the root radii approach with some advanced novel techniques, the paper [59] has noticeably accelerated the real root-finder [67], currently of user’s choice. One can extend the progress to fast approximation of near real roots by combining root radii computation with the techniques of Sec. 9.
(iii) Initialization based on Alg. D.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [65]. In particular Ehrlich’s iterations of MPSolve are celebrated for their very fast global empirical convergence (right from the start), but it has only been observed with such initialization; moreover, they diverge without it in an example of [125].

### v36

@@ 26.2 Definitions
$N(z):=\frac{p(z)}{p^{\prime}(z)},~{}W(x):=\frac{p(x)}{l^{\prime}(x)},~{}\frac{1}{E_{j}(x)}:=\frac{1}{N(x)}-\sum_{i=1,i\neq j}^{d}\frac{1}{x-z_{i}}$ (26.3)
@@
Newton’s iterations can be initiated at a single point $z$ or at many points and can approximate a single root, many or all $d$ roots at the cost decreasing as the number of roots decreases [71, 58, 143, 141, 85, 130].
@@
All expressions in (26.3) for the functions $N(x)$, $W(x)$, and $E_{j}(x)$ are already presented or can be rewritten as ratios $\frac{u(x)}{v(x)}$ for pairs of polynomials $u(x)$ and $v(x)$, whose coefficients are given or readily available at low cost if the polynomial $p(x)$ is gien with its coeffcients (see [102, Problems 3.6.1 and 3.6.3]). Then multipoint evaluation of such a function amounts to multipoint evaluation of $u(x)$ and $v(x)$ and division of the computed values, and one can accelerate Newton’s, Weierstrass’s, and Ehrlich’s iterations dramatically by incorporating the algorithms of [81, 59].
26.3 Acceleration with Fast Multipole Method (FMM)
The algorithms of [81, 59] deeply involve the coefficients of $p(x)$. Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding. Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [50] and by Carrier, Greengard, and Rokhlin [28] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The paper [121] proposes and extensively tests its acceleration to Superfast Multipole Method. The papers [105, 107] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
@@ 26.3 Acceleration with Fast Multipole Method (FMM)
$W(x)$ of (26.3) and all $j$ (cf. [16]). Let $I_{d}$ denote the $d\times d$ identity matrix, apply the Sherman-Morrison-Woodbury formula, and obtain
@@
$\frac{1}{p_{d}}p(x)=\det(xI_{d}-A)=-\sum_{i=1}^{d}w_{i}\prod_{j\neq i}(x-z_{j})+\prod_{j=1}^{d}(x-z_{j}).$ (26.4)
Indeed, the monic polynomials of degree $d$ on both sides have the same values for $x=z_{j}$, $j=1,\dots,d$ (cf. [16])). [^38] Gerschgorin’s theorem (cf. [48]) for the matrix $xI_{d}-A$ defines neighborhoods of its eigenvalues (the roots of $p$), which helps to analyze and to modify the iterations (cf. [22]). Equations (26.4) imply that
@@
$\frac{w_{i}}{N(z_{i})}=\sum_{j\neq i}\frac{w_{i}+w_{j}}{z_{i}-z_{j}}-1.$ (26.5)
Hence, given the $2d$ values of $z_{j}$ and $w_{j}=W(z_{j})$ for all $j$, one can fast approximate $N(z_{j})$ and consequently $E_{j}(z_{j})$ for all $j$ (cf. (26.3)) by applying FMM.
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (26.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (26.3) and (26.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [22, equation (17)]:
@@
(i) Use the following customary trick (see [32, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [32] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
@@
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 13.1. Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (3.3)) and can apply FMM to accelerate numerical computation of this sum. To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
26.4 MPSolve and its competition with Newton’s iterations
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [43], but the later version of MPSolve [22] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Schleicher and R. Stoll in [143] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see the previous subsection). The competition seems to resume in 2024 (see [12, 130, 85]).
26.5 Computation of a disc covering all zeros of a black box polynomial
26.5.1 Introductory comments
@@ 26.5.3 Deterministic solution
27 Fast approximation of the $d$ root radii with applications
The customary initialization recipes of Weierstrass"s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$. To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [10], reused in [15, 22], for the approximation of the $d$ root radii. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization rec;ipe turns out to be highly efficient.
The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s. It was first outlined and briefly analyzed in [134] for fast approximation of a single root radius and then extended in [92, Sec. 4] (cf. also [93, Sec. 5]) to approximation of all root radii of a polynomial. As in [10, 15, 22] it uses all $d$ coefficients of $p$. A variation of this algorithm in [62] supports the following theorem.
Theorem 27.1.
[62, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
@@ 27 Fast approximation of the $d$ root radii with applications
Combine Thm. 27.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [134, Cor. 14.3] and obtain
Corollary 27.1.
@@
The algorithm supporting the corollary and said to be Alg. 27.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [59], which extends the novel ingenious algorithm of [81] for multipoint polynomial evaluation.
@@
(i) In a subdivision step one can skip testing exclusion and discard a square unless it intersects the set $\mathbb{U}$.
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig. 11). [126] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments. According to these tests the algorithm is highly efficient for the approximation of all isolated real roots. One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec. 9. By complementing the root radii approach with some advanced novel techniques, the paper [62] has noticeably accelerated the real root-finder [70], which is currently user’s choice.
(iii) Initialization based on Alg. 27.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [68].